OpenStudy (anonymous):

Between which two integers does square root of 26 lie? A. 3 and 4 B. 5 and 6 C. 12 and 14 D. 25 and 27

OpenStudy (anonymous):

do you know how to do this @LegendarySadist

OpenStudy (anonymous):

one sec ill be back

OpenStudy (anonymous):

@phi

OpenStudy (phi):

to answer, first list the perfect squares "near" 26 can you do that ?

OpenStudy (anonymous):

um yeah 3 and 8 i believe

OpenStudy (anonymous):

or 5

OpenStudy (anonymous):

2,7,and one other

OpenStudy (anonymous):

is that right @phi

OpenStudy (phi):

do you know how to list squares? 1*1 = 1 2*2= 4 3*3= 9 keep doing that until you get two numbers that "bracket" 26 (one below and the other above)

OpenStudy (anonymous):

so do i have to get one that equals 26

OpenStudy (phi):

keep listing the perfect squares

OpenStudy (anonymous):

1*1=1 2*2=4 3*3=9 4*4=16 5*5=25 6*6=36

OpenStudy (phi):

notice 5*5 is 25 (1 less than 26) 6*6 is 36 (much bigger than 26) it looks like the number we want is a little bit bigger than 5

OpenStudy (anonymous):

okay so like 7*7=49

OpenStudy (anonymous):

oh do we need a smaller one than 26 to or were not that far yet

OpenStudy (phi):

you want a number times itself gives 26 for example, 5.1 * 5.1 = 26.01 which is very close the square root of 26 is very close to 5.1

OpenStudy (anonymous):

okay so maybe like 5.0990

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=what+is+the+square+root+of+26 heres how i got that

OpenStudy (phi):

it is hard to find the exact square root of 26 but we know the square root of 25 (it is 5) is too small and the square root of 36 (=6) is too big so the square root of 26 should be between 5 and 6 yes, 5.099 is a good guess, it gives 25.9998 which is 0.0002 too small. in the old days before calculators, people played that game to figure out square roots

OpenStudy (anonymous):

5.0990*5.0990 = 25.998 but if you round it to the nearest whole number it would =26

OpenStudy (anonymous):

opps i forgot a 9 2.9998

OpenStudy (phi):

yes. It depends how accurately you need to know the number, and that depends on the problem you are trying to solve. sometimes in engineering you need to be very accurate and sometimes less so. Depends on the problem

OpenStudy (anonymous):

okay so do we know what the answer would be then cuz with what we did it might be B

OpenStudy (phi):

yes, the square root of 26 is between 5 and 6 (as you just showed, the square root of 26 is very close to 5.099 and 5.099 is bigger than 5 and smaller than 6)

OpenStudy (anonymous):

okay so it is b thanks for helping me

OpenStudy (phi):

yw